OpenStudy (anonymous):

Find the slope of each line y=-2 y=x-2 y=1/2-4 3x+y=-5 6x-5y=5 x=-4 x-4y=0 4x-3y=-6

OpenStudy (anonymous):

wow.. use this man, y=mx +b where your m is the slope, and b as y-intercept... can you do that ?

OpenStudy (anonymous):

y=-2 slope is 0 y=x-2 slope is 1 y=1/2x-4 slope 1/2 3x+y=-5 slope is -3 6x-5y=5 slope is 6 x=-4 slope is undefined x-4y=0 slope is 1 4x-3y=-6 slope is 4

OpenStudy (anonymous):

oops.. I think you got it the wrong way.. this is how it works.. You must first get the equation into its proper form. For example, 3x+y=-5 y = -3x - 5, which is like y=mx+b so your m= -3, and b=-5 ^_^

OpenStudy (anonymous):

opps i should have solved it sorry but the slopes written r right thanks esgralitoculkin

OpenStudy (anonymous):

yeah... your welcome. ^_^